OpenStudy (anonymous):

5x2 – 6x – 9 = 0 Determine the discriminant for the equation. State the number of solutions. PLEASE HELP ME!!!

zepdrix (zepdrix):

Hey Rachel :) Recall that for a quadratic in standard form: \(\large\rm ax^2+bx+c=0\) Solutions are given by,\[\large\rm x=\frac{-b\pm\sqrt{\color{orangered}{b^2-4ac}}}{2a}\]This orange portion, under the root, is what we call the discriminant.

zepdrix (zepdrix):

If it turns out that the discriminant is negative, \(\large\rm \sqrt{\color{orangered}{-\text{#}}}\quad=\quad\sqrt{\text#}~i\) Then we end up with imaginary portion in our answer. So that corresponds to complex roots :)

zepdrix (zepdrix):

We really only care about the `sign` of the discriminant.

zepdrix (zepdrix):

a=5 b=-6 c=-9 Plug in the values, what do you get? :) b^2 - 4ac

OpenStudy (retireed):

Thanks zepdrix, I don't recall ever knowing that and now that I do, I can go to bed, since I learned something new today.

zepdrix (zepdrix):

Oh, and I guess since they're asking for "number of solutions", they're talking about "real solutions" here. So ignore what I said about complex roots :)

OpenStudy (anonymous):

okay so there would be 2 correct?

zepdrix (zepdrix):

Plugging in our values, it looks like we get...\[\large\rm b^2-4ac\quad=\quad (-6)^2-4(5)(-9)\quad=\quad 216\]Positive discriminant, so that corresponds to two roots, yes good job!

OpenStudy (anonymous):

Awesom!!! :)

zepdrix (zepdrix):

Lemme summarize that just in case you have more problems like this: If Discriminant = -# (no solutions) If Discriminant = 0 (one solution) If Discriminant = +# (two solutions)